Calculate the pH of a solution made by dissolving 50.0 mL of hydrogen chloride in water at 25 0C and 1.00 atm. The total volume of that solution is 455 mL.
1
Expert's answer
2020-04-09T13:11:40-0400

pH=-lg[H+]

[H+]=cHCl - because HCl is a strong acid in aqueous solutions, i.e. it dissociates into H+ and Cl- ions completely.

cHCl=nHCl/Vsolution

Vsolution=455mL=0.455L - the total volume of that solution is 455 mL

nHCl=50.0*10-3L / 22.4L*mol-1=0.00223mol, (22.4L*mol-1 - molar volume of gas under STD: 25 0C and 1.00 atm.)

[H+]=cHCl=0.00223mol / 0.455L=0.005mol/L

pH=-lg[H+]=-lg(0.005)=2.3


Answer: pH=2.3